# Fragmentation in Collapsar Disks: Migration, Growth, and Emission

### v1

@@ 3.3 Fragment Migration and Evolution
Figure 7: Fragment mass growth and inward migration time, plotted for each dimensionless radius $r/r_{\rm s}$ that the fragment passes through. We explore fragment evolution for a wide variety of different disk parameters $M$, $\dot{m}=\dot{M}/(M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$) and $\alpha$. Top panel: results for fixed $\alpha=0.02$ and $\dot{M}=1M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$, with multiple values of central mass $M$ shown as different evolutionary tracks. Middle panel: results for fixed $\alpha=0.02$ and $M=20M_{\odot}$, with multiple values of $\dot{m}$ shown as different evolutionary tracks. Here we also label each track by $\rho_{\star}\equiv\log_{10}(\rho)$, where $\rho$ is the initial fragment density in cgs normalized by ${\rm g\leavevmode\nobreak\ cm}^{-3}$. Bottom panel: results for fixed $M=20M_{\odot}$ and $\dot{M}=3M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$, with multiple values of $\alpha$ shown as different evolutionary tracks (that again are also labeled by $\rho_{\star}$). Color coding indicates the time it takes a fragment to reach a specific radius going inward from its initial formation radius on the right end of each track (denoted by a magenta ring, which corresponds to the magenta points in Fig. 7). Grey-scale points indicates failed assumptions, where: (1) light grey represents where the relative radial velocity is no longer negligible compared to the azimuthal relative velocity; (2) grey represents gap formation; (3) black represents when both a gap is formed and the relative radial velocity isn’t negligible. Fragments form with a typical density range of $10^{9}-10^{11}\leavevmode\nobreak\ {\rm gcm}^{-3}$ (above WD and below NS densities) and an initial mass of $\sim 10^{-3}M_{\odot}$, and grow by a few orders of magnitude to $\sim 1M_{\odot}$ where they form a gap in the disk and eventually migrate to the BH.
Fig. 7 shows three panels containing many evolutionary profiles of fragments experiencing migration and mass accumulation, where in each panel we vary either $M$, $\dot{m}$ or $\alpha$. We see that fragments form within a density range of $10^{9}-10^{11}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$, intermediate between white dwarf and neutron star interiors.
@@ 4 Observable Predictions
A major uncertainty in understanding the evolution of fragmentation arises from the initial masses and densities of self-bound fragments. Fragment initial masses almost always fall below the minimum stable cold neutron star mass (see Figs. 5, 6), and the initial densities generally fall into a range intermediate between white dwarf and neutron star densities (namely $\sim 10^{8-11}\leavevmode\nobreak\ {\rm g/cm}^{3}$). These objects cannot achieve a stable hydrostatic equilibrium in vacuum, even accounting for finite temperature effects (Gondek et al., 1997). In the absence of interaction with external media, the final fate of such objects is a delayed explosion (Blinnikov et al., 1990; Sumiyoshi et al., 1998). However, such an explosion is known to be delayed for dozens of seconds (Sumiyoshi et al., 1998), and in the complex environment of a collapsar disk, it may be additionally stabilized by the external ram pressure. [^9] When the fragment is still in its linear growth stage, ambient pressure from the collapsar disk will be important as well. The fragment may accrete enough mass to pass the minimum mass threshold before it explodes, or alternatively, migration to the ISCO may occur faster than the delayed explosion.
@@ 4.1 GRB Variability
Previous works (Perna et al., 2006; Dall’Osso et al., 2017; Shahamat & Abbassi, 2020; Shahamat et al., 2021) have considered the tidal disruption of migrating fragments as a potential source of GRB variability. Following disruption, a fragment returns its mass to the disk at a relatively small radius, creating a sizable overdensity in $\Sigma$ that then evolves viscously, temporarily enhancing the accretion rate onto the BH and modulating its jet power.
@@ 4.1.1 Secondary jet formation
Figure 9: Bondi-Hoyle accretion rates onto migrating fragments, $\dot{m}_{\rm f}$, normalized by the collapsar accretion rate $\dot{M}$ and plotted as functions of time since fragmentation. The color scheme shows instantaneous orbital radius $r/r_{\rm S}$). Top panel: results for fixed $\alpha=0.02$ and $\dot{M}=1M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$, with multiple values of central mass $M$ shown as different evolutionary tracks. Middle panel: results for fixed $\alpha=0.02$ and $M=20M_{\odot}$, with multiple values of $\dot{M}$ shown as different evolutionary tracks. Bottom panel: results for fixed $M=20M_{\odot}$ and $\dot{M}=3M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$, with multiple values of $\alpha$ shown as different evolutionary tracks. Grey-scale points indicates failed assumptions, with the same color-coding as in Fig. 7. The ratio $\dot{m}_{\rm f}/\dot{M}$ changes by multiple orders of magnitude during the fragment inspiral, rising from initial values $\sim 10^{-3}$ to final values $\sim 10^{1}$ (which are achieved briefly near the ISCO before terminating as the fragments plunge). At peak, $\dot{m}_{\rm f}$ generally exceeds $\dot{M}$, suggesting that secondary jets launched by the fragment may achieve significant luminosities (see 4.1.1). This plot does not account for fragment tidal disruption or the possibility of accretion rate suppression due to the neutrino Eddington limit.
@@ 4.2 Gravitational Waves
Here $\Theta$ is a dimensionless factor describing orientation of the source with respect to the antenna (spanning $0-4$; hereafter we choose $\Theta=2.5$, taking the mean square expectation of $\Theta$; Finn & Chernoff 1993), $m_{\rm GW}$ is the fragment mass at ${\rm max}(r_{{\rm t}},r_{{\rm Isco}})$ where $m_{{\rm 0.4}}\equiv m_{\rm GW}/0.4{M}_{\odot}$, the distance from the observer is set to be $100\leavevmode\nobreak\ {\rm Mpc}$, defining $D_{{\rm 100}}\equiv D_{{\rm obs}}/100{\rm Mpc}$, $M_{{\rm 20}}\equiv M/20{M}_{\odot}$ and $f_{{\rm 780}}\equiv f_{\rm GW}/780{\rm Hz}$ (as is appropriate for the ISCO frequency of the $a=0.95$ case).
@@
Figs. 11 and 12 show the GW strain and angular frequency, respectively, at the smallest radius we expect the fragment to survive at (i.e. the maximum of the tidal disruption radius, $r_{\rm t}$, and the ISCO). The maximum angular frequency is strongly correlated with the central mass because of the ISCO scaling. As the fragment migrates and grows with mass, it sweeps through an range of increasing angular frequencies. Because fragments usually form at large distances, their initial inspiral will not depend strongly on the central BH spin, but the maximum achievable frequency can be a sensitive function of spin (particularly for the largest fragments, which do not tidally disrupt). The range in maximum frequencies that we see in Fig. 11, where $a=0.95$, are in between $354-837\text{Hz}$. In contrast, the $a=0.001$ alternate case we have considered sees maximum frequencies between $88-350\leavevmode\nobreak\ {\rm Hz}$ (Fig. 26).
The maximum strain is likewise a sensitive function of spin for high mass fragments that avoid tidal disruption. As our scaling relation in Eq. 31c indicates, for a fiducial source at $D=100\leavevmode\nobreak\ {\rm Mpc}$, peak strains $h\sim 10^{-22}$ are achieved across most of the parameter space we consider. However, it is important to note that only some parameters allow us to estimate GW properties during the fragment inspiral. As discussed previously, fragmentation does not even occur for lower accretion rates ($\dot{M}\lesssim 0.4M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$) and higher effective viscosities ($\alpha\gtrsim 0.133$). For many of the smallest BHs and largest accretion rates we consider, we are unable to estimate GW emission, usually because of uncertainties about tidal disruption at large radii where the mass-radius relationship for hot, low mass fragments is unclear. For the smallest values of BH mass ($M\lesssim 10M_{\odot}$), we sometimes encounter other problems, such as fragments that never exceed the minimum cold neutron star mass, or numerical non-convergence of our disk model.
@@
[…]
˙$\dot{\Re}_{\rm LIGO}=1.08\cdot\chi_{\rm choke}f_{\rm f}N_{\rm f}\,\text{yr}^{-1}$ (32)
@@ 5 Discussion
To estimate plausible ranges of $E_{\gamma}/L_{\gamma}$ we use the long GRB catalogs of Tsvetkova et al. (2017, 2021), which compile 62 long GRBs with measured jet breaks and thus estimates for beaming-corrected $E_{\gamma}$ and $L_{\gamma}$. In this sample, typical values for the beaming-corrected energy release are $E_{\gamma}\sim 10^{49-52}\leavevmode\nobreak\ {\rm erg}$ and typical values for the beaming-corrected peak $\gamma$-ray luminosity are $L_{\gamma}\sim 10^{49-51.5}\leavevmode\nobreak\ {\rm erg\leavevmode\nobreak\ s}^{-1}$. The mean and median values of $t_{\rm peak}=E_{\gamma}/L_{\gamma}$ are 4.5 and 3.2 s, respectively; the 10th percentile and 90th percentile values for $t_{\rm peak}$ are 0.6 and 9.0 s, respectively. This implies that typical values of $\dot{M}\sim 1M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$ will be achieved if $\Delta M\sim 1M_{\odot}$; larger values will be reached if $\Delta M\sim 10M_{\odot}$. It is only if $\Delta M\ll 1M_{\odot}$ that the accretion rates required for fragmentation will not be reached in collapsar disks. Since we are primarily interested in central BHs with masses $\sim 10M_{\odot}$ that have grown by rapidly accreting a large fraction of their mass from the collapsing star’s core, high ($\gtrsim 1M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$) accretion rates are generally plausible unless the pre-collapse core was rotating so slowly that $\gtrsim 90\%$ of the BH mass was acquired through radial free-fall rather than through collapsar disk accretion.
@@ 5.1 Possible implications
While fragment disruption provides only modest GRB variability, a more significant effect may arise from the rapid accretion onto embedded fragments as they migrate inward. As fragments approach the ISCO, their accretion rates increase by several orders of magnitude, sometimes exceeding the central accretion rate. This could power secondary jets via neutrino annihilation or proto-magnetar spin-down, introducing variability between $15\%$ and $600\%$.
@@
If the fragmentation timescale turns out to be small compared to both the duration of the entire collapse event (the average time for LGRBs is in tens of seconds, Kouveliotou et al. 1993; Kumar & Zhang 2015; Bissaldi et al. 2017) and the fragment migration times, we hypothesize that fragments might collide and scatter D’Orazio & Loeb (2018) because of constant stochastic creation and rapid migration at different radii. For example if a fragment formed at an outer radius and starts migrating inwards it could overtake fragments that just finished forming at lower radii. collisions and mergers of high-density and high mass (that contracted and accreted significantly) fragmented bodies in a gas-rich environment, among themselves or with the BH, could produce non-vacuum GWs with electromagnetic counterparts (Piro & Pfahl, 2007; Fedrow et al., 2017; Metzger et al., 2024), while adding to GRB variability. Also, the presence of multiple bound fragments will lead to few-body scatterings (Stone & Leigh, 2019) that can eject single fragments or even binary fragments from the system. Since the end product of fragmentation with these extreme initial densities and fast mass growth is unclear, these ejected fragments could become compact objects such as BHs and NSs. An ejection of a compact binary might let them merge in a less gaseous environment, which would help produce the observed vacuum chirp signature seen by GW detectors.
@@
We have also estimated a significant GW detection rate from fragmentation in collapsars (As specified in Eq. 32 and under certain ranges of the highly unconstrained value of $\chi_{\rm choke}f_{\rm f}N_{\rm f}$). The derived detection rate ˙$\dot{\Re}_{\rm LIGO}=1.08\cdot\chi_{\rm choke}f_{\rm f}N_{\rm f},\text{yr}^{-1}$ suggests that fragment mergers within choked collapsars could substantially contribute to a new GW signal population. This highlights the potential of LIGO and future detectors to probe the dynamics of core collapse beyond traditional gamma-ray bursts.

### v2

@@ 3.3 Fragment Migration and Evolution
Figure 7: Fragment mass growth and inward migration time, plotted for each dimensionless radius $r/r_{\rm s}$ that the fragment passes through. We explore fragment evolution for a wide variety of different disk parameters $M$, $\dot{m}=\dot{M}/(M_{\odot}\penalty 10000\ {\rm s}^{-1}$) and $\alpha$. Top panel: results for fixed $\alpha=0.02$ and $\dot{M}=1M_{\odot}\penalty 10000\ {\rm s}^{-1}$, with multiple values of central mass $M$ shown as different evolutionary tracks. Middle panel: results for fixed $\alpha=0.02$ and $M=20M_{\odot}$, with multiple values of $\dot{m}$ shown as different evolutionary tracks. Here we also label each track by $\rho_{\star}\equiv\log_{10}(\rho)$, where $\rho$ is the initial fragment density in cgs normalized by ${\rm g\penalty 10000\ cm}^{-3}$. Bottom panel: results for fixed $M=20M_{\odot}$ and $\dot{M}=3M_{\odot}\penalty 10000\ {\rm s}^{-1}$, with multiple values of $\alpha$ shown as different evolutionary tracks (that again are also labeled by $\rho_{\star}$). Color coding indicates the time it takes a fragment to reach a specific radius going inward from its initial formation radius on the right end of each track (denoted by a magenta ring, which corresponds to the magenta points in Fig. 7). Grey-scale points indicates failed assumptions, where: (1) light grey represents where the relative radial velocity is no longer negligible compared to the azimuthal relative velocity; (2) grey represents gap formation; (3) black represents when both a gap is formed and the relative radial velocity isn’t negligible. Fragments form with a typical density range of $10^{9}-10^{11}\penalty 10000\ {\rm gcm}^{-3}$ (above WD and below NS densities) and an initial mass of $\sim 10^{-3}M_{\odot}$, and grow by a few orders of magnitude to $\sim 1M_{\odot}$ where they form a gap in the disk and eventually migrate to the BH.
Fig. 7 shows three panels containing many evolutionary profiles of fragments experiencing migration and mass accumulation, where in each panel we vary either $M$, $\dot{m}$ or $\alpha$. We see that fragments form within a density range of $10^{9}-10^{11}\penalty 10000\ {\rm g\penalty 10000\ cm}^{-3}$, intermediate between white dwarf and neutron star interiors.
@@ 4 Observable Predictions
A major uncertainty in understanding the evolution of fragmentation arises from the initial masses and densities of self-bound fragments. Fragment initial masses almost always fall below the minimum stable cold neutron star mass (see Figs. 5, 6), and the initial densities generally fall into a range intermediate between white dwarf and neutron star densities (namely $\sim 10^{8-11}\penalty 10000\ {\rm g/cm}^{3}$). These objects cannot achieve a stable hydrostatic equilibrium in vacuum, even accounting for finite temperature effects (Gondek et al., 1997). In the absence of interaction with external media, the final fate of such objects is a delayed explosion (Blinnikov et al., 1990; Sumiyoshi et al., 1998), determined by the timescale of the $\beta$-decays in those environments (Colpi et al., 1989; Sumiyoshi et al., 1998). However, such an explosion is known to be delayed for dozens of seconds (Sumiyoshi et al., 1998), and in the complex environment of a collapsar disk, it may be additionally stabilized by the external ram pressure. [^10] When the fragment is still in its linear growth stage, ambient pressure from the collapsar disk will be important as well. The fragment may accrete enough mass to pass the minimum mass threshold before it explodes, or alternatively, migration to the ISCO may occur faster than the delayed explosion (Chen & Metzger 2025 likewise noted that $\beta$-decay timescales could exceed the system’s evolution timescales, leading to extended survivability.).
@@ 4.1 GRB Variability
Previous works (Perna et al., 2006; Dall’Osso et al., 2017; Shahamat & Abbassi, 2020; Shahamat et al., 2021) have considered the tidal disruption of migrating fragments as a potential source of GRB variability. Following disruption, a fragment returns its mass to the disk at a relatively small radius, creating a sizable overdensity in $\Sigma$ that then evolves viscously, temporarily enhancing the accretion rate onto the BH and modulating its jet power. The typical prompt-emission variability timescales for long GRBs span from $\sim 0.1{\rm s}$, for the fastest spikes, up to tens of seconds for the slowest trends (Magnus et al., 2025).
@@ 4.1.1 Secondary jet formation
Figure 9: Bondi-Hoyle accretion rates onto migrating fragments, $\dot{m}_{\rm f}$, normalized by the collapsar accretion rate $\dot{M}$ and plotted as functions of time since fragmentation. The color scheme shows instantaneous orbital radius $r/r_{\rm S}$). Top panel: results for fixed $\alpha=0.02$ and $\dot{M}=1M_{\odot}\penalty 10000\ {\rm s}^{-1}$, with multiple values of central mass $M$ shown as different evolutionary tracks. Middle panel: results for fixed $\alpha=0.02$ and $M=20M_{\odot}$, with multiple values of $\dot{M}$ shown as different evolutionary tracks. Bottom panel: results for fixed $M=20M_{\odot}$ and $\dot{M}=3M_{\odot}\penalty 10000\ {\rm s}^{-1}$, with multiple values of $\alpha$ shown as different evolutionary tracks. Grey-scale points indicates failed assumptions, with the same color-coding as in Fig. 7. The ratio $\dot{m}_{\rm f}/\dot{M}$ changes by multiple orders of magnitude during the fragment inspiral, rising from initial values $\sim 10^{-3}$ to final values $\sim 10^{1}$ (which are achieved briefly near the ISCO before terminating as the fragments plunge). At peak, $\dot{m}_{\rm f}$ generally exceeds $\dot{M}$, suggesting that secondary jets launched by the fragment may achieve significant luminosities (see 4.1.1). This plot does not account for fragment tidal disruption or the possibility of accretion rate suppression due to the neutrino Eddington limit.
@@ 4.2 Gravitational Waves
Here $\Theta$ is a dimensionless factor describing orientation of the source with respect to the antenna (spanning 040-4; hereafter we choose $\Theta=2.5$, taking the mean square expectation of $\Theta$; Finn & Chernoff 1993), $m_{\rm GW}$ is the fragment mass at ${\rm max}(r_{{\rm t}},r_{{\rm Isco}})$ where $m_{{\rm 0.4}}\equiv m_{\rm GW}/0.4{M}_{\odot}$, the distance from the observer is set to be $100\penalty 10000\ {\rm Mpc}$, defining $D_{{\rm 100}}\equiv D_{{\rm obs}}/100{\rm Mpc}$, $M_{{\rm 20}}\equiv M/20{M}_{\odot}$ and $f_{{\rm 780}}\equiv f_{\rm GW}/780{\rm Hz}$ (as is appropriate for the ISCO frequency of the $a=0.95$ case).
@@
Figs. 11 and 12 show the GW strain and angular frequency, respectively, at the smallest radius we expect the fragment to survive at (i.e. the maximum of the tidal disruption radius, $r_{\rm t}$, and the ISCO). The maximum angular frequency is strongly correlated with the central mass because of the ISCO scaling. As the fragment migrates and grows with mass, it sweeps through an range of increasing angular frequencies. Because fragments usually form at large distances, their initial inspiral will not depend strongly on the central BH spin, but the maximum achievable frequency can be a sensitive function of spin (particularly for the largest fragments, which do not tidally disrupt). The range in maximum frequencies that we see in Fig. 11, where $a=0.95$, are in between $354-837\text{Hz}$. In contrast, the $a=0.001$ alternate case we have considered sees maximum frequencies between $88-350\penalty 10000\ {\rm Hz}$ (Fig. 26).
The maximum strain is likewise a sensitive function of spin for high mass fragments that avoid tidal disruption. As our scaling relation in Eq. 31c indicates, for a fiducial source at $D=100\penalty 10000\ {\rm Mpc}$, peak strains $h\sim 10^{-22}$ are achieved across most of the parameter space we consider. However, it is important to note that only some parameters allow us to estimate GW properties during the fragment inspiral. As discussed previously, fragmentation does not even occur for lower accretion rates ($\dot{M}\lesssim 0.4M_{\odot}\penalty 10000\ {\rm s}^{-1}$) and higher effective viscosities ($\alpha\gtrsim 0.133$). For many of the smallest BHs and largest accretion rates we consider, we are unable to estimate GW emission, usually because of uncertainties about tidal disruption at large radii where the mass-radius relationship for hot, low mass fragments is unclear. For the smallest values of BH mass ($M\lesssim 10M_{\odot}$), we sometimes encounter other problems, such as fragments that never exceed the minimum cold neutron star mass, or numerical non-convergence of our disk model.
@@
We have estimated a typical maximum strain of $\sim 1.5\cdot 10^{-22}$. Taking a LIGO $\rm O4$} strain limit for detection [^13] We estimate this limit crudely by taking a detection horizon of $160$ Mpc for the merger of two $1.4M_{\odot}$ NSs, each with a radius of $10$ km and $\Theta=2.5$. of $\sim 5\cdot 10^{-23}$, the corresponding maximum detection distance for fragment-BH mergers is approximately $295\penalty 10000\ {\rm Mpc}$. From these estimates, we derive the LIGO detection rate ($\dot{\Re}_{\rm LIGO}$) of such GW signals as a function of $\chi_{\rm choke}$, the fraction of collapsars that generate fragments $f_{\rm f}\leq 1$, and the mean number of fragments that merge with the BH $N_{\rm f}\geq 1$:
$\dot{\Re}_{\rm LIGO}=1.08\cdot\chi_{\rm choke}f_{\rm f}N_{\rm f}\,\text{yr}^{-1}$ (32)
@@ 5 Discussion
To estimate plausible ranges of $E_{\gamma}/L_{\gamma}$ we use the long GRB catalogs of Tsvetkova et al. (2017, 2021), which compile 62 long GRBs with measured jet breaks and thus estimates for beaming-corrected $E_{\gamma}$ and $L_{\gamma}$. In this sample, typical values for the beaming-corrected energy release are $E_{\gamma}\sim 10^{49-52}\penalty 10000\ {\rm erg}$ and typical values for the beaming-corrected peak $\gamma$-ray luminosity are $L_{\gamma}\sim 10^{49-51.5}\penalty 10000\ {\rm erg\penalty 10000\ s}^{-1}$. The mean and median values of $t_{\rm peak}=E_{\gamma}/L_{\gamma}$ are 4.5 and 3.2 s, respectively; the 10th percentile and 90th percentile values for $t_{\rm peak}$ are 0.6 and 9.0 s, respectively. This implies that typical values of $\dot{M}\sim 1M_{\odot}\penalty 10000\ {\rm s}^{-1}$ will be achieved if $\Delta M\sim 1M_{\odot}$; larger values will be reached if $\Delta M\sim 10M_{\odot}$. It is only if $\Delta M\ll 1M_{\odot}$ that the accretion rates required for fragmentation will not be reached in collapsar disks. Since we are primarily interested in central BHs with masses $\sim 10M_{\odot}$ that have grown by rapidly accreting a large fraction of their mass from the collapsing star’s core, high ($\gtrsim 1M_{\odot}\penalty 10000\ {\rm s}^{-1}$) accretion rates are generally plausible unless the pre-collapse core was rotating so slowly that $\gtrsim 90\%$ of the BH mass was acquired through radial free-fall rather than through collapsar disk accretion.
@@ 5.1 Possible implications
While fragment disruption provides only modest GRB variability, a more significant effect may arise from the rapid accretion onto embedded fragments as they migrate inward. As fragments approach the ISCO, their accretion rates increase by several orders of magnitude, sometimes exceeding the central accretion rate. This could power secondary jets via neutrino annihilation or proto-magnetar spin-down, introducing variability between $15\%$ and $600\%$. The secondary jets appear to exhibit two distinct timescales: the migration time of the fragment until it reaches gap formation—halting accretion and secondary jet power—which is roughly in the range of $\sim 0.1$$1{\rm s}$ (see Fig. 9); and the duration during which the accretion rate of the fragment exceeds the central accretion rate (see Fig. 28), which is $\sim 1$$10{\rm ms}$. The migration timescale could plausibly be related to the variability timescales observed in long GRBs, where the fastest spikes in the prompt emission typically occur on timescales that are greater than $\sim 0.1{\rm s}$ (Magnus et al., 2025).
@@
If the fragmentation timescale turns out to be small compared to both the duration of the entire collapse event (the average time for LGRBs is in tens of seconds, Kouveliotou et al. 1993; Kumar & Zhang 2015; Bissaldi et al. 2017) and the fragment migration times, we hypothesize that fragments might collide and scatter (D’Orazio & Loeb, 2018) because of constant stochastic creation and rapid migration at different radii. For example, if a fragment formed at an outer radius migrates inwards, it could overtake fragments that formed at lower radii. Close encounters can lead to binary formation via gas dissipation (Li et al., 2023b), in turn producing mergers of high-density and high mass (following accretion fragments in a gas-rich environment. This exciting possibility could produce non-vacuum GWs with electromagnetic counterparts (Piro & Pfahl, 2007; Fedrow et al., 2017; Metzger et al., 2024), while also adding to GRB variability. Also, the presence of multiple bound fragments will lead to few-body scatterings (Stone & Leigh, 2019) that can eject single fragments or even binary fragments from the system. Since the end product of fragmentation with these extreme initial densities and fast mass growth is unclear, these ejected fragments could become compact objects such as BHs and NSs. An ejection of a compact binary might let them merge in a less gaseous environment, which would help produce the observed vacuum chirp signature seen by GW detectors.
@@
We have also estimated a significant GW detection rate from fragmentation in collapsars (As specified in Eq. 32 and under certain ranges of the highly unconstrained value of $\chi_{\rm choke}f_{\rm f}N_{\rm f}$). The derived detection rate $\dot{\Re}_{\rm LIGO}=1.08\cdot\chi_{\rm choke}f_{\rm f}N_{\rm f},\text{yr}^{-1}$ suggests that fragment mergers within choked collapsars could substantially contribute to a new GW signal population. This highlights the potential of LIGO and future detectors to probe the dynamics of core collapse beyond traditional gamma-ray bursts.